OpenStudy (anonymous):

Can somebody explain how my online math course got from (x-y)y/(y-x)x to -y/x?

OpenStudy (llort):

\[xy-y^2/xy-x^2\]

OpenStudy (anonymous):

?

OpenStudy (somethingawesome):

Notice that -(x-y) = -x+y = y-x. So, to swap the order like that, we can factor out a -1. Then \[ \frac{(x-y)y}{(y-x)x} = \frac{-(y-x)y}{(y-x)x} = \frac{-y}{x}\] after we cancel the (y-x) on the top and bottom.

OpenStudy (anonymous):

i'm confused..

OpenStudy (anonymous):

did'nt you just multiply the numerator by negative 1?

OpenStudy (anonymous):

oh, wait..I got it now...

OpenStudy (anonymous):

how do I give you a medal again?

OpenStudy (somethingawesome):

Glad to help! I think you just click that "Good Answer" button.

OpenStudy (anonymous):

I don't see the 'good answer' button..

OpenStudy (somethingawesome):

It's usually next to everyone else's post, on the right. Don't worry about it if it's not there.

OpenStudy (llort):

I gave it for you...

OpenStudy (somethingawesome):

Thanks :)